OpenStudy (kmullis6):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). a) y = 1 over 3x + 7 b) y = -3x + 3 c) y = 1 over 3x + 3 d) y = -3x + 7

OpenStudy (anonymous):

parallel lines have the same slope so we know the slope of the equation of the line that is parallel will be -3

OpenStudy (anonymous):

IF it is parallel to a line with a slope of -3 like this one, that means it will also have a slope of -3. THis means that a and c are out. Now we will use point slope from \(y-y1=m(x-x1)\) where m=-3, x1=-1, and y1=6. can you plug those in?

OpenStudy (anonymous):

 I know it sucks. well plug that in with the slope and tell me what you get